# The radiative decay of scalar glueball from lattice QCD

### v2

@@ Abstract
We perform the first lattice QCD study on the radiative decay of the scalar glueball to the vector meson $\phi$ in the quenched approximation. The calculations are carried out on three gauge ensembles with different lattice spacings, which enable us to do the continuum extrapolation. We first revisit the radiative $J/\psi$ decay into the scalar glueball $G$ and obtain the partial decay width $\Gamma(J/\psi\to\gamma G)=0.449(44)~{}\text{keV}$ and the branching fraction $\text{Br}(J/\psi\to\gamma G)=4.8(5)\times 10^{-3}$, which are in agreement with the previous lattice results. We then extend the similar calculation to the process $G\to\gamma\phi$ and get the partial decay width $\Gamma(G\to\gamma\phi)=0.074(47)~{}\text{keV}$, which implies that the combined branching fraction of $J/\psi\to\gamma G\to\gamma\gamma\phi$ is as small as $\mathcal{O}(10^{-9})$ such that this process is hardly detected by the BESIII experiment even with the large $J/\psi$ sample of $\mathcal{O}(10^{10})$. With the vector meson dominance model, the two-photon decay width of the scalar glueball is estimated to be $\Gamma(G\to\gamma\gamma)=0.53(46)~{}\text{eV}$, which results in a large stickiness $S(G)\sim\mathcal{O}(10^{4})$ of the scalar glueball by assuming the stickiness of $f_{2}(1270)$ to be one.
@@ I INTRODUCTION
Gluons and quarks are fundamental degrees of freedom of the Quantum Chromodynamics (QCD). Apart from the conventional mesons and baryons that are described by quark-antiquark ($q\bar{q}$) and three quark ($qqq$) bound states in the constituent quark models, it is usually conjectured that there also exist glueballs that are bound states of pure gluons. Glueballs are well-defined objects in the pure Yang-Mills theory, whose spectrum have been derived through the numerical lattice QCD calculations in the quenched approximation [1, 2]. For example, the lowest lying scalar ($0^{++}$), tensor ($2^{++}$) and pseudoscalar ($0^{-+}$) glueball masses are predicted to be $1.5-1.7$ GeV, $2.2-2.4$ GeV and $2.4-2.6$ GeV, respectively. These results are supported to some extent by recent dynamical lattice QCD [3, 4, 5, 6]. Obviously, these lowest lying glueballs share the same quantum numbers with the conventional $q\bar{q}$ mesons, and can mix with $q\bar{q}$ mesons when the gluon-quark transition is switch on. The key question is to single out the (predominant) glueball states among mesons of the same quantum numbers and similar masses.
@@ III.2 Two-point function
Figure 1: The effective mass of $\phi$ correlators. From top to bottom, the results correspond to $\beta=2.4$, $2.8$, $3.0$, respectively, from bottom to top are the plateaus for momentum modes, $\vec{p}=\frac{2\pi\vec{n}}{La_{s}}$, with $|\vec{n}|^{2}=0,1,2,...,9$.
Figure 2: Dispersion relationship of $\phi$ mesons on three ensembles. The results from top to bottom are for $\beta=2.4$, $\beta=2.8$, and $\beta=3.0$. Two methods are used to verify the dispersion relationship. When assuming the speed of light is equal to 1 and fitting the dispersion relationship by $E_{G}^{2}(\vec{p})a_{t}^{2}=m_{G}^{2}a_{t}^{2}+\frac{1}{\xi^{2}}|\vec{p}|^{2}a_{s}^{2}$, the resulting anisotropy $\xi$ values are 5.07, 5.04, and 5.06, respectively, as shown in the larger graphs. Alternatively, using a bare anisotropy of $\xi=5$, the speed of light is calculated by $c^{2}=\frac{E^{2}(\vec{p})-m_{\phi}^{2}}{|\vec{p}|^{2}}$, as depicted in the smaller graphs.
In order to determine the masses of the $\phi$ meson and glueball, it is necessary to calculate the corresponding two-point function. The local fermion bilinear operators, denoted as $\mathcal{O}=\psi\Gamma\bar{\psi}$, are utilized to compute the mass spectra of strangeonium, as outlined in Tab. 1. To enhance the signal-to-noise ratio, point sources are placed on each time slice to calculate propagators, and subsequently, the resulting two-point functions are averaged.
@@
We employ the method for constructing scalar glueball operators as outlined in reference [28]. By constructing 24 glueball operators set $\{\mathcal{O}_{G,i},i=1,2,...,24\}$ comprised of Wilson loops satisfying the $A_{1}^{++}$ representation of the discrete spatial point group, we assemble the two-point function matrix $\Gamma_{G,ij}^{(2)}(t,\vec{p})=\langle\mathcal{O}_{G,i}(t,\vec{p})\mathcal{O}^{\dagger}_{G,j}(0)\rangle$. By performing a variational analysis on this $24\times 24$ correlation function matrix and solving the generalized eigenvalue equation, we obtain optimized operators $\mathcal{O}_{G,\lambda}=\sum_{i}c_{\lambda,i}\mathcal{O}_{G,i}$ that project predominantly onto single states, where $c_{\lambda,i}$ are the eigenvectors corresponding to the eigenvalue $\lambda$ of the generalized eigenvalue equation. Given the substantial statistical fluctuations of the glueball operators, the construction of optimized glueball operators is crucial for studying the glueball spectrum and its decay properties. We performed a single mass term fit to the optimized glueball two-point function normalized using $\Gamma_{G}^{(2)}(t=0,\vec{p})$, and found that the overlap factor $Z_{G}$ is very close to 1. This indicates that the optimized glueball operator almost projects onto the ground scalar glueball. Furthermore, given the short effective fitting range for the scalar glueball, we also considered the systematic errors arising from different choices of fitting intervals. The effective mass results from the optimized glueball operators are shown in Fig. 3 and the glueball mass obtained from our fitting is presented in Tab. 1. These optimized glueball operators will be further utilized for constructing the three-point functions discussed later.
@@ III.3 $J/\psi\to\gamma G$
Figure 4: Extrapolation of $E_{1}(Q^{2})$ to the on-shell point $Q^{2}=0$ for the $J/\psi\to\gamma G$ process. The results from top to bottom correspond to $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$.
@@
Here, we place the $J/\psi$ operator $\mathcal{O}_{J/\psi}$ at $t_{i}=0$, the current operator at time $t$, and the optimized scalar glueball operator at time $t_{f}$. Due to our use of the optimized glueball operator, we were able to set $t_{f}$ very close to $t$. In fact, we tested three scenarios with $\Delta t\equiv t_{f}-t=0,1,2$, and found that the results did not show a significant dependence on $\Delta t$. In subsequent calculations, we present the results for $\Delta t=1$ and consider the impact of different $\Delta t$ as a source of systematic error. The difference from our previous work [28] is that we use a wall source when computing the quark propagator, which reduces the accessible values of $Q^{2}$, but provides a better signal-to-noise ratio. We define the ratio
@@
where $\Gamma_{J/\psi}^{(2)}$, $\Gamma_{G}^{(2)}$ are $J/\psi$ and scalar glueball two-point correlation. The overlap factors $W_{J/\psi}(\vec{p}_{i})=\frac{Z_{J/\psi}^{2}}{2E_{J/\psi}V}$, $W_{G}(\vec{p}_{f})=\frac{Z_{G}^{2}}{2E_{G}V}$, as well as the energy of $J/\psi$ $E_{J/\psi}(\vec{p}_{i})$ and the scalar glueball $E_{G}(\vec{p}_{f})$ can be obtained by fitting from corresponding the two-point functions. It should be noted that, the local vector current $J_{\text{em}}^{\mu}=\bar{c}\gamma^{\mu}c$ here and $J_{\text{em}}^{\mu}=\bar{s}\gamma^{\mu}s$ for the process $G\to\gamma\phi$, which are conserved in the continuum, are no longer conserved on the finite lattice and requires multiplicative normalization factors $Z_{V}^{\mu}$, which can be determined following the strategy in Refs. [38, 46]. Only the spatial components of $J_{\text{em}}^{\mu}$ are involved in our calculation, whose normalization factors ($Z_{V}^{(c)}$ for $\bar{c}\gamma_{i}c$ and $Z_{V}^{(s)}$ for $\bar{s}\gamma_{i}s$) are collected in Tab. 2, where the values of $Z_{V}^{(c)}$ are taken from Ref. [47], while those of $Z_{V}^{(s)}$ are derived in this work. In practice, these normalization factors are incorporated implicitly in the calculation.
@@
$Z_{V}^{(c)}$ $1.388(15)$ $1.110(7)$ $1.062(6)$
$Z_{V}^{(s)}$ $0.903(22)$ $0.955(16)$ $1.008(30)$
@@
Our fitting is shown in Fig. 4. By using wall source to compute the charm quark propagator, we effectively improve the signal of three point function, resulting in a smoother variation of the form factor $E_{1}(Q^{2})$ with $Q^{2}$ compared to the literature [28]. After obtaining $E_{1}(0)$ for each ensemble, we further extrapolate the results to the continuum limit using a linear function:
@@
$E_{1}(0)^{\text{cont. }}=0.0595(29)~{}\text{GeV},$ (11)
@@
$\Gamma_{J/\psi\rightarrow\gamma G}=0.449(44)~{}\text{keV}.$ (12)
This result is consistent with the previous lattice QCD result $\Gamma_{J/\psi\to\gamma G}=0.35(8)~{}\text{keV}$ [28].
@@ III.4 $G\to\gamma\phi$
Based on the Eq. 3, for the calculation of the process $G\to\gamma\phi$, we need to place the glueball operator at the source, that is, at the time slice $t_{i}=0$, and then multiply it by the two-point correlation function of the current and $\phi$ operators. However, in practical calculations, we find that the three-point function obtained in this way has very large statistical noise, making it difficult to obtain an effective signal. Considering the crossing symmetry of the matrix element
$\langle\phi(\vec{p}_{f})|J_{\text{em}}^{\mu}(x)|G(\vec{p}_{i})\rangle^{*}=\langle G(-\vec{p}_{i})|J_{\text{em}}^{\mu}(x)|\phi(-\vec{p}_{f})\rangle,$ (13)
we can also construct the three-point function of this process similarly to the $J/\psi\to\gamma G$ process, that is,
$\Gamma_{\phi\gamma\to G}^{(3),\mu}(t_{f},t)=\langle\mathcal{O}_{G}(\vec{p}_{f},t_{f})J_{\text{em}}^{s,\mu}(\vec{q},t)\mathcal{O}_{\phi}^{\dagger}(\vec{p}_{i},0)\rangle.$ (14)
@@
Figure 8: The variation of the form factors $E_{1}(Q^{2})$ with the time interval $\Delta t=t_{f}-t$ is presented. The cases for $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$ are shown sequentially from top to bottom. For the cases of $\beta=2.8$ and $\beta=3.0$, only three time intervals $\Delta t=0,1,2$ were investigated. In all cases, the form factors exhibit a linear dependence on the time interval.
@@ IV DISCUSSION
where $V_{\mu},G$ are the fields for the vector meson ($J/\psi$ or $\phi$ in this work) and the scalar glueball, respectively, $A_{\mu}$ is the electromagnetic field, $-e$ is the electric charge of electron, $Q_{q}$ is the electric charge of the quark in the vector meson, and $g_{\text{eff}}$ describes the $V\leftrightarrow G$ transition. Considering the expressions in Eq. (2), one can see $g_{\text{eff}}=E_{1}(0)$. If we introduce a dimensionless coupling constant $g_{G\phi}=E_{1}(0)/m_{G}$ for $G\to\gamma\phi$ and a coupling constant $g_{G\psi}=E_{1}(0)/m_{\psi}$ for $J/\psi\to\gamma G$, then we get
$g_{G\phi}/g_{G\psi}=0.73(41)\sim\mathcal{O}(1)$ (33)
@@
The partial decay width of $J/\psi\to\gamma G$ is predicted to be $0.449(44)~{}\text{keV}$, which gives the branching fraction
$\text{Br}(J/\psi\to\gamma G)=4.8(5)\times 10^{-3},$ (34)
if the total width of $J/\psi$ $\Gamma=92.6(1.7)~{}\text{keV}$ [24] is used. These results are in agreement with the previous lattice results $0.35(8)~{}\text{keV}$ and $3.8(9)\times 10^{-3}$ [28]. As for the scalar glueball candidates $f_{0}(1710)$ and $f_{0}(1500)$, summing up the PDG data of the branching fractions of $J/\psi\to\gamma f_{0}(1710)\to\gamma(K\bar{K},\pi\pi,\eta\eta,\omega\omega,\omega\phi)$ [24] gives the lower bound $\mathrm{Br}(J/\psi\to\gamma f_{0}(1710))>2.1\times 10^{-3}$, while the lower bound for $f_{0}(1500)$ is $\mathrm{Br}(J/\psi\to\gamma f_{0}(1500))>1.9\times 10^{-4}$ by summing up the branching fractions of $J/\psi\to\gamma f_{0}(1500)\to\gamma(\pi\pi,\eta\eta,K\bar{K})$. Obviously, the production rate of $f_{0}(1710)$ in the $J/\psi$ radiative decay is one order of magnitude larger than that of $f_{0}(1500)$ are consistent with the production rate of the scalar glueball. On the other hand, BESII and BESIII have performed the partial wave analysis (PWA) of the processes $J/\psi\to\gamma X\to\gamma\pi\pi$ [25], $\gamma\eta\eta$ [26], $\gamma K_{S}K_{S}$ [27], and found that in each process, $f_{0}(1710)$ is produced much more than $f_{0}(1500)$. These observations support $f_{0}(1710)$ to be predominantly a scalar glueball state or have a large component of the scalar glueball. Through a coupled channel analysis to $J/\psi\to\gamma(\pi\pi,K\bar{K},\eta\eta,\phi\omega)$ processes and considering the octet-singlet mixings of scalar mesons, Klempt et al. claim that there should be a glueball state with the parameters $(m,\Gamma)=(1865\pm 25_{-30}^{+10},370\pm 50_{-20}^{+30})~{}\text{MeV}$, and its observed yield in radiative $J/\psi$ decays is $5.8(1.0)\times 10^{-3}$ [53, 54].
@@
where $C\approx 17.7$ is taken to make the stickiness of $f_{2}(1270)$ to be $S(f_{2}(1270))=1$ using the PDG data [24], $q_{\gamma}=(m_{\psi}^{2}-m_{G}^{2})/(2m_{\psi})$ is the momentum of the photon in the process $J/\psi\to\gamma G$ (in the rest frame of $J/\psi$). This value can be use as a reference to identify a pure glueball state or estimate the glueball component of a meson in the future experiments. The phenomenological study based on the non-relativistic gluon bound state model also predicts the partial width $\Gamma(G\to\gamma\gamma)$ to have similar magnitude [55].

### v3

@@ Abstract
We perform the first lattice QCD study on the radiative decay of the scalar glueball to the vector meson $\phi$ in the quenched approximation. The calculations are carried out on three gauge ensembles with different lattice spacings, which enable us to do the continuum extrapolation. We first revisit the radiative $J/\psi$ decay into the scalar glueball $G$ and obtain the partial decay width $\Gamma(J/\psi\to\gamma G)=0.578(86)~{}\text{keV}$ and the branching fraction $\text{Br}(J/\psi\to\gamma G)=6.2(9)\times 10^{-3}$. We then extend the similar calculation to the process $G\to\gamma\phi$ and get the partial decay width $\Gamma(G\to\gamma\phi)=0.074(47)~{}\text{keV}$, which implies that the combined branching fraction of $J/\psi\to\gamma G\to\gamma\gamma\phi$ is as small as $\mathcal{O}(10^{-9})$ such that this process is hardly detected by the BESIII experiment even with the large $J/\psi$ sample of $\mathcal{O}(10^{10})$. With the vector meson dominance model, the two-photon decay width of the scalar glueball is estimated to be $\Gamma(G\to\gamma\gamma)=0.53(46)~{}\text{eV}$, which results in a large stickiness $S(G)\sim\mathcal{O}(10^{4})$ of the scalar glueball by assuming the stickiness of $f_{2}(1270)$ to be one.
pacs:
11.15.Ha, 12.38.Gc, 12.39.Mk, 13.30.Ce
@@ I INTRODUCTION
Gluons and quarks are fundamental degrees of freedom of the Quantum Chromodynamics (QCD). Apart from the conventional mesons and baryons that are described by quark-antiquark ($q\bar{q}$) and three quark ($qqq$) bound states in the constituent quark models, it is usually conjectured that there also exist glueballs that are bound states of pure gluons. Glueballs are well-defined objects in the pure Yang-Mills theory, whose spectrum have been derived through the numerical lattice QCD calculations in the quenched approximation [1, 2, 3]. For example, the lowest lying scalar ($0^{++}$), tensor ($2^{++}$) and pseudoscalar ($0^{-+}$) glueball masses are predicted to be $1.5-1.7$ GeV, $2.2-2.4$ GeV and $2.4-2.6$ GeV, respectively. These results are supported to some extent by recent dynamical lattice QCD [4, 5, 6, 7, 8]. Obviously, these lowest lying glueballs share the same quantum numbers with the conventional $q\bar{q}$ mesons, and can mix with $q\bar{q}$ mesons when the gluon-quark transition is switch on. The key question is to single out the (predominant) glueball states among mesons of the same quantum numbers and similar masses.
@@ III.2 Two-point function
Figure 1: The effective mass of $\phi$ correlators. From top to bottom, the results correspond to $\beta=2.4$, $2.8$, $3.0$, respectively, from bottom to top are the plateaus for momentum modes, $\vec{p}=\frac{2\pi\vec{n}}{La_{s}}$, with $|\vec{n}|^{2}=0,1,2,...,9$. The band represent the range of two-mass fitting.
Figure 2: Dispersion relationship of $\phi$ mesons on three ensembles. The results from top to bottom are for $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$. Two methods are used to verify the dispersion relationship. When assuming the speed of light is equal to 1 and fitting the dispersion relationship by $E_{\phi}^{2}(\vec{p})a_{t}^{2}=m_{\phi}^{2}a_{t}^{2}+\frac{1}{\xi^{2}}|\vec{p}|^{2}a_{s}^{2}$, the resulting anisotropy $\xi$ values are 5.08, 5.02, and 5.00, respectively, as shown in the larger graphs. Alternatively, using a bare anisotropy of $\xi=5$, the speed of light is calculated by $c^{2}=\frac{E^{2}(\vec{p})-m_{\phi}^{2}}{|\vec{p}|^{2}}$, as depicted in the smaller graphs.
Figure 3: Dispersion relationship of glueball on three ensembles. The results from top to bottom are for $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$.
In order to determine the masses of the $\phi$ meson and glueball, it is necessary to calculate the corresponding two-point function. The local fermion bilinear operators, denoted as $\mathcal{O}=\bar{\psi}\Gamma\psi$, are utilized to compute the mass spectra of strangeonium, as outlined in Tab. 1. To enhance the signal-to-noise ratio, point sources are placed on each time slice to calculate propagators, and subsequently, the resulting two-point functions are averaged.
@@
We employ the method for constructing scalar glueball operators as outlined in reference [30]. By constructing 24 glueball operators set $\{\mathcal{O}_{G,i},i=1,2,...,24\}$ comprised of Wilson loops satisfying the $A_{1}^{++}$ representation of the discrete spatial point group, we assemble the two-point function matrix $\Gamma_{G,ij}^{(2)}(t,\vec{p})=\langle\mathcal{O}_{G,i}(t,\vec{p})\mathcal{O}^{\dagger}_{G,j}(0)\rangle$. By performing a variational analysis on this $24\times 24$ correlation function matrix and solving the generalized eigenvalue equation, we obtain optimized operators $\mathcal{O}_{G,\lambda}=\sum_{i}c_{\lambda,i}\mathcal{O}_{G,i}$ that project predominantly onto single states, where $c_{\lambda,i}$ are the eigenvectors corresponding to the eigenvalue $\lambda$ of the generalized eigenvalue equation. Given the substantial statistical fluctuations of the glueball operators, the construction of optimized glueball operators is crucial for studying the glueball spectrum and its decay properties. We performed a single mass term fit to the optimized glueball two-point function normalized using $\Gamma_{G}^{(2)}(t=0,\vec{p})$, and found that the overlap factor $Z_{G}$ is very close to 1. This indicates that the optimized glueball operator almost projects onto the ground scalar glueball. Furthermore, given the short effective fitting range for the scalar glueball, we also considered the systematic errors arising from different choices of fitting intervals. The dispersion relationship and effective mass results from the optimized glueball operators are shown in Fig. 3 and Fig. 4 respectively, and the glueball mass obtained from our fitting is presented in Tab. 1. These optimized glueball operators will be further utilized for constructing the three-point functions discussed later.
@@ III.3 $J/\psi\to\gamma G$
Figure 5: Extrapolation of $E_{1}(Q^{2})$ to the on-shell point $Q^{2}=0$ for the $J/\psi\to\gamma G$ process. The results from left to right correspond to $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$.
@@
Here, we place the $J/\psi$ operator $\mathcal{O}_{J/\psi}$ at $t_{i}=0$, the current operator at time $t$, and the optimized scalar glueball operator at time $t_{f}$. Due to our use of the optimized glueball operator, we were able to set $t_{f}$ very close to $t$. In fact, we tested three scenarios with $\Delta t\equiv t_{f}-t=0,1,2$, and found that the results did not show a significant dependence on $\Delta t$. In subsequent calculations, we use the results with $\Delta t=2$ for $\beta=2.4$ and $\beta=2.8$, and the result with $\Delta t=5$ for $\beta=3.0$. The difference from our previous work [30] is that we use a wall source when computing the quark propagator, which reduces the accessible values of $Q^{2}$, but provides a better signal-to-noise ratio. We define the ratio
@@
where $\Gamma_{J/\psi}^{(2)}$, $\Gamma_{G}^{(2)}$ are $J/\psi$ and scalar glueball two-point correlation. The overlap factors $W_{J/\psi}(\vec{p}_{i})=\frac{Z_{J/\psi}^{2}}{2E_{J/\psi}V}$, $W_{G}(\vec{p}_{f})=\frac{Z_{G}^{2}}{2E_{G}V}$, as well as the energy of $J/\psi$ $E_{J/\psi}(\vec{p}_{i})$ and the scalar glueball $E_{G}(\vec{p}_{f})$ can be obtained by fitting from corresponding the two-point functions. It should be noted that, the local vector current $J_{\text{em}}^{\mu}=\bar{c}\gamma^{\mu}c$ here and $J_{\text{em}}^{\mu}=\bar{s}\gamma^{\mu}s$ for the process $G\to\gamma\phi$, which are conserved in the continuum, are no longer conserved on the finite lattice and requires multiplicative normalization factors $Z_{V}^{\mu}$, which can be determined following the strategy in Refs. [40, 48]
$Z_{V_{f=s,c}}^{(\mu)}(t)=\frac{p^{\mu}}{E(\vec{p})}\frac{\Gamma_{\eta_{f}}^{(2)}\left(\vec{p};t_{f}\right)}{\Gamma_{\eta_{f}}^{(3),\mu}\left(\vec{p},\vec{p},t_{f},t\right)}.$ (8)
Only the spatial components $J_{\text{em}}^{i=x,y,z}$ are involved in our calculation, whose renormalization factors ($Z_{V_{c}}^{(i)}$ for $\bar{c}\gamma_{i}c$ and $Z_{V_{s}}^{(i)}$ for $\bar{s}\gamma_{i}s$) are collected in Tab. 2
@@
$Z_{V_{c}}^{(i)}$ [^3] The values of $Z^{(i)}_{V_{c}}$ listed in Refs. [48, 30, 49] include an additional factor $\nu_{s}$ to suppress the discrepancy between $Z_{V_{c}}^{(i)}$ and $Z_{V_{c}}^{(t)}$, which would cancel part of the discretization errors, especially on two coarse lattices with $\beta=2.4$ and 2.8 [48, 30]. When we included one more fine lattice $(\beta=3.0)$ to the calculation of the decay process $J/\psi\to\gamma G_{0^{-+}}$ in Ref. [49], we realized the $\nu_{s}$ factor had been inappropriately added and was removed in the practical calculation. The values of $Z_{V_{c}}^{(i)}$ in Table 2 are the ones after removing the $\nu_{s}=0.71,0.74,0.77$ at $\beta=2.4,2.8,3.0$, respectively. The values from Refs. [48, 50] should be updated accordingly. $1.955(21)$ $1.500(10)$ $1.379(8)$
@@
Our fitting is shown in Fig. 5. Considering the applicability of polynomial interpolation, we present the interpolation results for data points within $Q^{2}\in[-2,2]~{}\text{GeV}^{2}$. By using wall source to compute the charm quark propagator, we effectively improve the signal of three point function, resulting in a smoother variation of the form factor $E_{1}(Q^{2})$ with $Q^{2}$ compared to the literature [30]. After obtaining $E_{1}(0)$ for each ensemble, we further extrapolate the results to the continuum limit using a linear function:
@@
$E_{1}(0)^{\text{cont. }}=0.0675(50)~{}\text{GeV},$ (12)
@@
$\Gamma_{J/\psi\rightarrow\gamma G}=0.578(86)~{}\text{keV}.$ (13)
Due to the change in current renormalization, our result is slightly larger than the previous result, $\Gamma_{J/\psi\to\gamma G}=0.35(8)~{}\text{keV}$ [30], but it is consistent within $2\sigma$.
@@ III.4 $G\to\gamma\phi$
Based on the Eq. 3, for the calculation of the process $G\to\gamma\phi$, we need to place the glueball operator at the source, that is, at the time slice $t_{i}=0$, and then multiply it by the two-point correlation function of the current and $\phi$ operators
$\Gamma_{G\to\gamma\phi}^{(3),\mu}(t_{f},\vec{p}_{f};t,\vec{q})$ $=\langle\mathcal{O}_{\phi}(\vec{p}_{f},t_{f})J_{em}^{s,\mu}(\vec{q},t)\mathcal{O}_{G}^{\dagger}(\vec{p}_{i},0)\rangle$
$=\sum_{x,y,z}e^{-i\vec{p}_{f}z}e^{i\vec{q}(x+y)}\langle\bar{\psi}(z,t_{f})\gamma_{i}\psi(z,t_{f})$
×$\times\bar{\psi}(x,t)\gamma_{\mu}\psi(y,t)\mathcal{O}_{G}^{\dagger}(\vec{p}_{i},0)\rangle.$ (14)
However, in practical calculations, we find that the three-point function obtained in this way has very large statistical noise, making it difficult to obtain an effective signal. we can also construct the three-point function of this process similarly to the $J/\psi\to\gamma G$ process, that is,
$\Gamma_{\phi\gamma\to G}^{(3),\mu}(t_{f},\vec{p}_{f};t,\vec{q})$ $=\langle\mathcal{O}_{G}(\vec{p}_{f},t_{f})J_{em}^{s,\mu}(\vec{q},t)\mathcal{O}_{\phi}(\vec{p}_{i},0)\rangle$
$=\sum_{x,y,z}e^{i\vec{q}z}e^{i\vec{p}_{i}(x+y)}\langle\mathcal{O}_{G}^{\dagger}(\vec{p}_{f},t_{f})$
×$\times\bar{\psi}$ $(z,t)\gamma_{\mu}\psi(z,t)\bar{\psi}(x,0)\gamma_{i}\psi(y,0)\rangle.$ (15)
@@
Figure 7: The variation of the form factors $E_{1}(Q^{2})$ with the time interval $\Delta t=t_{f}-t$ is presented. The cases for $\beta=2.4$, $\beta=2.8$ and $\beta=3.0$ are shown sequentially from left to right. For the cases of $\beta=2.8$ and $\beta=3.0$, only three time intervals $\Delta t=0,1,2$ were investigated. In all cases, the form factors exhibit a linear dependence on the time interval.
@@ IV DISCUSSION
Glueballs, as bound states of gluons, are well defined objects in pure Yang-Mills theories. Their existence has been confirmed by previous lattice QCD calculations in the quenched approximation [1, 2, 3]. There are also some lattice QCD calculations with dynamical quarks using the gluonic operators to predict the glueball masses and having obtained results consistent with those from the quenched approximation [4, 5, 6, 7, 8]. Strictly speaking, the definition of glueballs in full QCD is not so conceptually rigorous as that in the quenched approximation owing to the gluon-quark transition. With this in mind, the phenomenological studies of hadrons usually conjecture that glueballs do exist (at least as a degree of freedom) and a physical meson state can be a $q\bar{q}$ meson, a multiquark state, a glueball, or an admixture of them. In this sense, the glueball properties derived in the quenched appproximation serve as theoretical inputs to the phenomenological interpretation of the nature of a meson, especially to identify its glueball components. This is the motivation of our discussion in the following.
@@
where $V_{\mu},G$ are the fields for the vector meson ($J/\psi$ or $\phi$ in this work) and the scalar glueball, respectively, $A_{\mu}$ is the electromagnetic field, $-e$ is the electric charge of electron, $Q_{q}$ is the electric charge of the quark in the vector meson, and $g_{\text{eff}}$ describes the $V\leftrightarrow G$ transition. Considering the expressions in Eq. (2), one can see $g_{\text{eff}}=E_{1}(0)$. If we introduce a dimensionless coupling constant $g_{G\phi}=E_{1}(0)/m_{G}$ for $G\to\gamma\phi$ and a coupling constant $g_{GJ/\psi}=E_{1}(0)/m_{J/\psi}$ for $J/\psi\to\gamma G$, then we get
$g_{G\phi}/g_{GJ/\psi}=0.73(41)\sim\mathcal{O}(1)$ (34)
@@
The partial decay width of $J/\psi\to\gamma G$ is predicted to be $0.578(86)~{}\text{keV}$, which gives the branching fraction
$\text{Br}(J/\psi\to\gamma G)=6.2(9)\times 10^{-3},$ (35)
when the $J/\psi$ total width $\Gamma=92.6(1.7)~{}\text{keV}$ [26] is used. These results are slightly larger than but qualitatively compatible with the previous lattice results $0.35(8)~{}\text{keV}$ and $3.8(9)\times 10^{-3}$ [30]. As for the scalar glueball candidates $f_{0}(1710)$ and $f_{0}(1500)$, summing up the PDG data of the branching fractions of $J/\psi\to\gamma f_{0}(1710)\to\gamma(K\bar{K},\pi\pi,\eta\eta,\omega\omega,\omega\phi)$ [26] gives the lower bound $\mathrm{Br}(J/\psi\to\gamma f_{0}(1710))>2.1\times 10^{-3}$, while the lower bound for $f_{0}(1500)$ is $\mathrm{Br}(J/\psi\to\gamma f_{0}(1500))>1.9\times 10^{-4}$ by summing up the branching fractions of $J/\psi\to\gamma f_{0}(1500)\to\gamma(\pi\pi,\eta\eta,K\bar{K})$. Obviously, the production rate of $f_{0}(1710)$ in the $J/\psi$ radiative decay is one order of magnitude larger than that of $f_{0}(1500)$ are consistent with the production rate of the scalar glueball. On the other hand, BESII and BESIII have performed the partial wave analysis (PWA) of the processes $J/\psi\to\gamma X\to\gamma\pi\pi$ [27], $\gamma\eta\eta$ [28], $\gamma K_{S}K_{S}$ [29], and found that in each process, $f_{0}(1710)$ is produced much more than $f_{0}(1500)$. These observations support $f_{0}(1710)$ to be predominantly a scalar glueball state or have a large component of the scalar glueball. Through a coupled channel analysis to $J/\psi\to\gamma(\pi\pi,K\bar{K},\eta\eta,\phi\omega)$ processes and considering the octet-singlet mixings of scalar mesons, Klempt et al. claim that there should be a glueball state with the parameters $(m,\Gamma)=(1865\pm 25_{-30}^{+10},370\pm 50_{-20}^{+30})~{}\text{MeV}$, and its observed yield in radiative $J/\psi$ decays is $5.8(1.0)\times 10^{-3}$ [56, 57, 58].
@@
where $C\approx 17.7$ is taken to make the stickiness of $f_{2}(1270)$ to be $S(f_{2}(1270))=1$ using the PDG data [26], $q_{\gamma}=(m_{J/\psi}^{2}-m_{G}^{2})/(2m_{J/\psi})$ is the momentum of the photon in the process $J/\psi\to\gamma G$ (in the rest frame of $J/\psi$). This value can be use as a reference to identify a pure glueball state or estimate the glueball component of a meson in the future experiments. The phenomenological study based on the non-relativistic gluon bound state model also predicts the partial width $\Gamma(G\to\gamma\gamma)$ to have similar magnitude [59].